OpenStudy (anonymous):

sin 2θ cos θ − cos 2θ sin θ =root2/2

OpenStudy (anonymous):

the right side becomes sin θ by the difference identity for sine so θ = pi/4 or 3pi/4.

OpenStudy (anonymous):

nope its wrong

OpenStudy (dumbcow):

sin 2x = 2sinx cosx cos 2x = 2cos^2 -1 --> 2sinx cos^2 -2sinx cos^2 + sinx = root2/2